40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6724     Entry time: Thu May 31 01:27:16 2012
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: PSL and Y arm green beams aligned 

[Jenne, Yuta]

We aligned the PSL green optics so that the PSL green beam and Y arm green beam interfere. 2 beams are now hitting the Y arm beat PD. The DC level from the beat PD is about 13 mV.

We didn't try to see the beat signal for today, because the temperature of the doubling crystal seemed funny. We need to look into it tommorow.

Currently, the temperature control is enabled and the set point is 36.9 deg C, but the temperature is stuck at 33.0 deg C.

ELOG V3.1.3-